Firm X : sells goods to firm Y to produce its service of value of 200€, sells good to final consumers at 500€, pay wages of 200€

Firm Y : sells goods to firm X to produce its service of value of 150€, sells good to final consumers at 200€, it sells abroad for 100, pay wages of 150€


a) Estimate GDP using production and income methods.
b) Assume that Firm X decides to transfer its production abroad. Would the GDP of the country under analysis increase, decrease or remain the same and why ?
c) Assume that Firm Y decides to continue its production in the country but that foreign multinatinal corporation decides to buy it. Would the GDP of the country under analysis increase, decrease or remain the same and why ?
1
Expert's answer
2021-02-03T16:10:01-0500

Solution:

a.). Estimate GDP using production and income methods.

GDP using production method:

Gross value Output at market price - value of Intermediate Consumption = Gross value addedMP (GDP)

Value added Good X = GVOMP – IC = GVAMP

                                                   = 500 – 200 = 300

Value added Good Y = 200 – 150 = 50

GDP                                               = 350


GDP using income method:

 = Total National Income + Taxes + Depreciation + Net foreign factor income

= 200 + 150 = 350

GDP = 350


b.). The GDP of the country will decrease. This is because all factors of production are added to attain GDP and by removing them, the GDP will decrease. GDP measures the total value of final goods and services produced within a given country’s borders.


c.). The GDP of the country will remain the same. This is because foreign ownership does not affect GDP values. Whether a company is locally owned or foreign-owned, what matters is that the production is done locally and goods are also consumed locally.